# Denoising as Adaptation: Noise-Space Domain Adaptation for Image Restoration

### v1

@@ 3.2 Eliminating Shortcut Learning in Diffusion
To avoid the above shortcut in diffusion, as shown in Fig. 4, we first propose a channel shuffling layer $f_{cs}$ to randomly shuffle the channel index of synthetic and real-world conditions at each iteration before concatenating them, i.e., $\mathbf{C}(f_{cs}(\hat{\bm{y}}^{s},\hat{\bm{y}}^{r}))$ [^2] We omit the shuffling operator $f_{cs}$ for notation clarity in the following presentation.. We show in the experiments that this strategy is crucial to bridge the gap of synthetic and real data.
In addition to channel shuffling, we also devise residual-swapping contrastive learning to ensure the network learns to restore genuinely instead of overfitting the paired synthetic appearance. Using the ground truth noise $\bm{\epsilon}$ as the anchor, we construct a positive example $\bm{\epsilon}^{pos}$ derived from Eq. (1): $\bm{\epsilon}^{pos}=\bm{\epsilon}_{\theta}\left(\tilde{\bm{y}}^{s},\mathbf{C}(\hat{\bm{y}}^{s},\hat{\bm{y}}^{r}),t\right)$, i.e., the expected noise from the diffusion model conditioning on restored synthetic and real-world images. We then swap the residual maps of these two conditions and formulate a negative example $\bm{\epsilon}^{neg}$ as follows:
$\bm{\epsilon}^{neg}$ $=\bm{\epsilon}_{\theta}\left(\tilde{\bm{y}}^{s},\mathbf{C}(\hat{\bm{y}}^{s\leftarrow r},\hat{\bm{y}}^{r\leftarrow s}),t\right),~{}\hat{\bm{y}}^{s\leftarrow r}=\bm{x}^{s}\oplus\mathcal{R}^{r},~{}\hat{\bm{y}}^{r\leftarrow s}=\bm{x}^{r}\oplus\mathcal{R}^{s},$ (2)
where $\mathcal{R}^{s}$ and $\mathcal{R}^{r}$ are the estimated residual maps of the corresponding synthetic image and real-world image from the restoration network, and $\oplus$ is the pixel-wise addition operator. By swapping the residual maps of two conditions, we constrain the diffusion model to repel the distance between the wrong restored results and the expected clean distribution regardless of their context. Based on the positive, negative, and anchor examples, the residual-swapping contrastive learning can be formulated as:
@@
where $\delta$ denotes a predefined margin to separate the positive and negative samples. In this way, the loss of diffusion model takes the mean of Eq. (1) and Eq. (3).
In the above formulation, the synthetic restored image of the condition, denoted as $\hat{\bm{y}}^{s}$, and the input to the diffusion model, represented as $\tilde{\bm{y}}^{s}$, form a pair of data with evident pixel-wise similarity. This similarity can potentially mislead the diffusion model to ignore the real restored image $\hat{\bm{y}}^{r}$ in condition as analyzed in Fig. 3. It is important to note that the distribution $\mathcal{S}$ encapsulates the domain knowledge of high-quality clean images, including but not limited to the ground truth images in the synthetic dataset. Motivated by this observation, the proposed method can be further extended by replacing the noisy input $\tilde{\bm{y}}^{s}$ with $\tilde{\bm{y}}^{c}$, defined as $\tilde{\bm{y}}^{c}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{c}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where ${\bm{y}}^{c}$ is randomly sampled from an unpaired extensive high-quality image dataset. This strategy disrupts the pixel-wise similarity between the “synthetic” condition and the diffusion input, thus enforcing the diffusion model to guide both the “synthetic” and “real” conditions predicted by the restoration network at the domain level. We will provide an ablation on this setting in Sec. 4.1.
@@ 3.3 Training
The image restoration network and diffusion model are jointly optimized by:
@@
Following previous works [19], we gradually change $\lambda_{Dif}$ from 00 to $\beta$ to avoid distractions for the main image restoration task during the early stages of the training process:
@@ 3.4 Discussion
The proposed denoising as adaption is reminiscent of the domain adversarial objective proposed by Ganin and Lempitsky [19]. The main difference is that we do not use a domain classifier with a gradient reversal layer but a diffusion network for the loss. We categorize methods like [19] as feature-space domain adaptation approaches. Unlike these approaches, we show that denoising as adaptation is more well-suited for image restoration as it can better preserve low-level appearance in the pixel-wise noise space. Compared to pixel-space approaches that usually require multiple generator and discriminator networks, our method adopts a compact framework incorporating only a single additional denoising U-Net, ensuring stable adaptation training. After training, the diffusion network is discarded, requiring only the learned restoration network for testing purposes. The framework comparison of the above three types of methods is presented in Sec. B of the Appendix.
@@ 4 Experiments
Training Dataset. For image denoising, we follow previous works [34, 31] and construct the synthetic training dataset based on DIV2K [72], Flickr2K [73], WED [74], and BSD [75]. The noisy images are obtained by adding the additive white Gaussian noise (AWGN) of noise level $\sigma\in[0,75]$ to the source clean images. We use the training dataset of SIDD [76] as the real-world data. For image deraining, the synthetic and real-world training datasets are respectively obtained from Rain13K [45] and SPA [77]. For image deblurring, GoPro [38] and RealBlur-J [78] are selected as the synthetic and real-world training datasets, respectively. Please note that we only use the degraded images from these real-world datasets (without the ground truth) for training purposes. For large-scale unpaired clean images, all images in the MS-COCO dataset [79] are used.
Testing Dataset. The testing images of the real-world datasets (SIDD [76], SPA [77], RealBlur-J [78]) are employed to evaluate the performance of the corresponding image restoration models.
Training Settings. To train the diffusion model, we adopt $\alpha$ conditioning and the linear noise schedule ranging from $-61e\text{-}$ 6 to $-21e\text{-}$ 2 following previous works [61, 62, 80]. Moreover, the EMA strategy with a decaying factor of $0.9999$ is also used across our experiments. Both the restoration and diffusion networks are trained on $128\times 128$ patches, which are processed with random cropping and rotation for data augmentation. Our model is trained with a fixed learning rate $-55e\text{-}$ 5 using Adam [81] algorithm and the batch size is set to 40. All experiments are conducted on NVIDIA A100 GPUs.
Metrics. The performance of various methods is mainly evaluated using the classical metrics: PSNR, SSIM, and LPIPS. For the image deraining task, we calculate PSNR/SSIM scores using the Y channel in YCbCr color space following existing methods [42, 43, 31].
4.1 Comparisons with State-of-the-Art Methods
We implement the proposed noise-space domain adaptation method using a handy and classical U-Net architecture [82]. To validate its effectiveness, we compare the proposed method with previous domain adaptation approaches, including DANN [19], DSN [22], PixelDA [27], and CyCADA [51], covering the feature-space and pixel-space adaptation solutions. For the purpose of a fair comparison, we retrained these methods with the same standard settings and datasets. In addition, we also consider some unsupervised image restoration methods and representative supervised methods such as Ne2Ne [14], MaskedD [15], NLCL [16], SelfDeblur [5], VDIP [17], and Restormer [31].
Figure 5: Visual comparison of the image denoising task on SIDD test dataset [76].
Comparison Results. The quantitative and qualitative comparison results are shown in Tab. 1-3 and Fig. 5-6. From the comparison results, the proposed method leads the comparison methods on three image restoration tasks. In particular, previous feature-space domain adaptation methods [19, 22, 51] fail to perceive the crucial low-level information and pixel-space domain adaptation methods [27, 51] yield inferior results since the precise style transfer between two domains is hard to control during the adversarial training. Moreover, the self-supervised and unsupervised restoration methods [14, 15, 5, 16, 17] show noticeable artifacts and limited generalization performance due to some inevitable information loss and hand-crafted designs on specific degradations. By contrast, our method ensures a fine domain adaptation in the pixel-wise noise space without introducing unstable training.
Table 1: Quantitative evaluation of the image denoising task on SIDD test dataset [76]. syn, real, both denote the model is trained on synthetic, real-world (w/o GT), and both synthetic and real-world (w/o GT) datasets, respectively. The best score is highlighted with color shading.
Metrics Vanilla DANN [19] DSN [22] PixelDA [27] CyCADA [51] Ne2Ne [14] MaskedD [15] Ours
@@ 4.1 Comparisons with State-of-the-Art Methods
Table 2: Quantitative evaluation of the image deraining task on SPA test dataset [77]. PSNR/SSIM scores are calculated using the Y channel in the YCbCr color space.
Metrics Vanilla DANN [19] DSN [22] PixelDA [27] CyCADA [51] NLCL [16] Restormer [31] Ours
@@
Table 3: Quantitative evaluation of the image deblurring task on RealBlur-J test dataset [78].
Metrics Vanilla DANN [19] DSN [22] PixelDA [27] CyCADA [51] SelfDeblur [5] VDIP [17] Ours
@@
Figure 6: Visual comparison of the image deraining and image deblurring tasks on SPA [77] and RealBlur-J [78] test datasets.
@@
Analysis. From the above results, we can observe that the proposed method enables noticeable improvements beyond the Vanilla baseline (trained only with synthetic datasets) on the image restoration tasks involved with high-frequency noises, such as image denoising and image deraining. Especially for image denoising, $+8.13/0.3070$ improvements on PSNR/SSIM metrics are achieved. We argue that the target of image denoising naturally fits that of the forward denoising process in the diffusion model. It is more sensitive to other Gaussian-like noises with respect to the pre-sampled noise space. Thus, an intense diffusion loss would be back-propagated if the conditioned images are under-restored, and the preceding restoration network tries to eliminate the noises on both the synthetic and real-world images as much as possible.
Table 4: Quantitative metrics of the proposed method (Ours) and its extension on unpaired condition case (Our-Ex). The results are formed with PSNR/SSIM/LPIPS. The best and second best scores are highlighted and underlined.
Task Ours Ours-Ex
Denoising 34.71/0.9202/0.0903 33.44/0.8938/0.1064
Deraining 34.39/0.9571/0.0462 34.20/0.9587/0.0444
Deblurring 26.46/0.8048/0.1363 26.44/0.8030/0.1313
Extension. As mentioned in Sec. 3.2, our method can extend to the unpaired condition case by relaxing the diffusion’s input with the image from other clean datasets. Thus, the shortcut issue can be directly eliminated since the trivial solutions such as matching the pixel’s similarity between input and condition do not exist. Such an extension keeps the channel shuffling layer but is free to the residual swapping contrastive learning. We show the quantitative evaluation in Tab. 4. The results demonstrate that although the condition and diffusion input are unpaired, our method can still learn to adapt the restored results from the synthetic and real-world domains to the clean image distribution, which also complements the restoration performance of the paired solution in some tasks. More qualitative results are presented in the Appendix.
@@
Figure 7: Visual comparison results of ablation studies.
4.2 Ablation Studies
To evaluate the effectiveness of different components in the proposed method, we conduct ablation studies regarding the sampled noise levels of the diffusion model, determined by the time-step $t$, and the training strategies to avoid shortcut learning, as shown in Tab. 5 and Fig. 7. Concretely, with low noise intensity, e.g., $t\in[1,100]$, it is easy for the diffusion model to discriminate the similarity of paired synthetic data even when the restored conditions are under-restored. As a result, the shortcut learning comes earlier during the training process and the real-world degraded image is heavily corrupted by the restoration network, of which most all details are filtered. On the other hand, when the intensity of the sampled noise is high, e.g., $t\in[900,1000]$, the diffusion model is hard to converge and the whole framework has fallen into a local optimum. By sampling the noise from a more diverse range with $t\in[1,1000]$, the restored results can be gradually adapted to the clean distribution. Moreover, the generalization ability of the restoration network gains further improvement using the designed channel shuffling layer (CS) and residual-swapping contrastive learning strategy (RS), which effectively eliminates the shortcut learning of the diffusion model. Therefore, higher restoration performance on real-world images and more realistic visual appearance can be observed from (d) to (e) and (f) in Tab. 5 and Fig. 7.
@@ 4.3 Scalability
We further validate the scalability of the proposed method, using different variants of U-Net-based image restoration networks and other types of architectures such as the Transformer-based network [32]. In particular, we classify these networks based on their model sizes and obtain: Unet-T, Unet-S (the model investigated in the above experiments), Unet-B, Uformer-T, Uformer-S, and Uformer-B. More network details are listed in the Appendix. The quantitative results of PSNR vs. computational cost on SIDD test dataset [76] are shown in Fig. 8. As we can observe, as the complexity and parameter increase, the vanilla restoration network (orange elements) tends to overfit the training synthetic dataset and perform worse on the test real-world dataset. In contrast, the proposed domain adaptation approach method can improve the generalization ability of image restoration models with various sizes and architectures (blue elements). It is also interesting that for each type of architecture, our method can facilitate better adaptation performance as the complexity of the restoration network increases, demonstrating its effectiveness in addressing the overfitting problem of large models.
4.4 Limitation and Broader Impacts
In the proposed noise-space domain adaptation, we discard the diffusion model after its joint training with the image restoration network. Like previous domain adaptation methods, we need to retrain the proxy model once new datasets or tasks are involved. However, our learned diffusion model has a strong capacity to discriminate if the restored results are good or inferior. Thus, it could be leveraged to directly provide the prior knowledge to inspire new restoration-related tasks. Consequently, the adaptation from the synthetic domain to a new real-world domain would be achieved more efficiently. We leave it as one of our future work. For the broader impacts of this work, image restoration can improve the quality and accessibility of visual information in various fields, including medical imaging, satellite imagery, and historical document preservation. It can benefit better diagnostic accuracy in healthcare, more precise environmental monitoring, and the safeguarding of cultural heritage. While the image restoration network can vastly improve degraded images, it also carries the risk of inadvertently editing and altering the original information, which may compromise the authenticity of the restored images.

### v2

@@ 3.2 Eliminating Shortcut Learning in Diffusion
To avoid the above shortcut in the diffusion model, as shown in Fig. 3, we first propose a channel shuffling layer $f_{cs}$ to randomly shuffle the channel index of synthetic and real-world conditions at each iteration before concatenating them, i.e., $\mathbf{C}(f_{cs}(\hat{\bm{y}}^{s},\hat{\bm{y}}^{r}))$ [^2] We omit the shuffling operator $f_{cs}$ for notation clarity in the following presentation.. We show in the experiments that this strategy is important to bridge the gap between synthetic and real data.
In addition to channel shuffling, we devise a residual-swapping contrastive learning strategy to ensure the network learns to restore genuinely instead of overfitting the paired synthetic appearance. Using the ground truth noise $\bm{\epsilon}$ as the anchor, we construct a positive example $\bm{\epsilon}^{pos}$ derived from Eq. 1: $\bm{\epsilon}^{pos}=\bm{\epsilon}_{\theta}\left(\tilde{\bm{y}}^{s}|\mathbf{C}(\hat{\bm{y}}^{s},\hat{\bm{y}}^{r}),t\right)$, i.e., the expected noise from the diffusion model conditioned on restored synthetic and real-world images. We then swap the residual maps of these two conditions and formulate a negative example $\bm{\epsilon}^{neg}$ as follows:
$\bm{\epsilon}^{neg}$ $=\bm{\epsilon}_{\theta}\left(\tilde{\bm{y}}^{s}|\mathbf{C}(\hat{\bm{y}}^{s\leftarrow r},\hat{\bm{y}}^{r\leftarrow s}),t\right),~{}\hat{\bm{y}}^{s\leftarrow r}=\bm{x}^{s}\oplus\mathcal{R}^{r},~{}\hat{\bm{y}}^{r\leftarrow s}=\bm{x}^{r}\oplus\mathcal{R}^{s},$ (2)
where $\mathcal{R}^{s}$ and $\mathcal{R}^{r}$ are the estimated residual maps of the synthetic and real-world images from the restoration network, and $\oplus$ is the pixel-wise addition operator. By swapping the residual of two conditions, we constrain the diffusion model to repel the distance between the wrong restored results and the expected clean distribution regardless of their context relation. Based on the positive, negative, and anchor examples, a compact residual-swapping contrastive learning can be formulated as:
@@
where $\delta$ denotes a predefined margin to separate the positive and negative samples. In this way, the loss of diffusion model takes the mean of Eq. 1 and Eq. 3. By implementing the above strategies, we challenge the diffusion model to distinguish between synthetic and real conditions based on trivial solutions, thereby encouraging both to align with the target clean distribution.
In the above formulation, the synthetic restored image of the condition, denoted as $\hat{\bm{y}}^{s}$, and the input to the diffusion model, represented as $\tilde{\bm{y}}^{s}$, form a pair of data with evident pixel-wise similarity. This similarity can potentially mislead the diffusion model to ignore the real restored image $\hat{\bm{y}}^{r}$ in condition as analyzed in Fig. 2. It is important to note that the target distribution encapsulates the domain knowledge of high-quality clean images, including but not limited to the ground truth images in the synthetic dataset. Motivated by this observation, the proposed method can be further extended by replacing the noisy input $\tilde{\bm{y}}^{s}$ with $\tilde{\bm{y}}^{c}$, defined as $\tilde{\bm{y}}^{c}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{c}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where ${\bm{y}}^{c}$ is randomly sampled from an unpaired extensive high-quality image dataset. This strategy disrupts the pixel-wise similarity between the synthetic condition and the diffusion input, thus enforcing the diffusion model to guide both the synthetic and real conditions predicted by the restoration network at the domain level. We will provide an ablation on this setting in Appendix A4.1.
@@ 3.3 Training
In the proposed training strategy, the restoration and diffusion models are jointly optimized by:
@@
Following previous works (Ganin & Lempitsky, 2015), we gradually change $\lambda_{Dif}$ from 00 to $\beta$ to avoid distractions for the main image restoration task during the early stages of the training process:
@@ 3.4 Discussion
The proposed denoising as adaption is reminiscent of the domain adversarial objective proposed by  (Ganin & Lempitsky, 2015). The main difference is that we do not use a domain classifier with a gradient reversal layer but a diffusion network for the loss. We categorize methods like (Ganin & Lempitsky, 2015) as feature-space domain adaptation approaches. Unlike these approaches, we show that denoising as adaptation is more well-suited for image restoration as it can better preserve low-level appearance in the pixel-wise noise space. Compared to pixel-space approaches that usually require multiple generator and discriminator networks, our method adopts a compact framework incorporating only a single additional denoising U-Net, ensuring stable adaptation training. After training, the diffusion network is discarded, requiring only the restoration network for inference. The framework comparison of the above three types of methods is presented in Appendix  A2.
@@ 4 Experiments
Dataset. For image denoising, we follow previous works (Zhang et al., 2018a; Zamir et al., 2022) and construct the synthetic training dataset based on DIV2K (Timofte et al., 2017), Flickr2K (Nah et al., 2019), WED (Ma et al., 2016), and BSD (Martin et al., 2001). The noisy images are obtained by adding the additive white Gaussian noise (AWGN) of noise level $\sigma\in[0,75]$ to the source clean images. We use the training dataset of SIDD (Abdelhamed et al., 2018) as the real-world data. For image deraining, the synthetic and real-world training datasets are respectively obtained from Rain13K (Yang et al., 2017) and SPA (Wang et al., 2019). For image deblurring, GoPro (Nah et al., 2017) and RealBlur-J (Rim et al., 2020) are selected as the synthetic and real-world training datasets, respectively. Please note that we only use the degraded images from these real-world datasets (without the ground truth) for training purposes. For large-scale unpaired clean images, all images in the MS-COCO dataset (Lin et al., 2014) are used. The test images of the real-world datasets (SIDD, SPA, RealBlur-J) are employed to evaluate the performance of the corresponding image restoration models.
Figure 4: Visual comparison of the image denoising task on SIDD test dataset (Abdelhamed et al., 2018).
Table 1: Quantitative evaluation of the image denoising task on SIDD test dataset (Abdelhamed et al., 2018). syn, real, both denote the model is trained on synthetic, real-world (w/o GT), and both synthetic and real-world (w/o GT) datasets, respectively. The best score is highlighted.
Metrics Vanilla DANN DSN PixelDA CyCADA Ne2Ne MaskedD Ours
@@
Training Settings. To train the diffusion model, we adopt $\alpha$ conditioning and the linear noise schedule ranging from $-61e\text{-}$ 6 to $-21e\text{-}$ 2 following previous works (Saharia et al., 2022a; c; Chen et al., 2020). Moreover, the EMA strategy with a decaying factor of $0.9999$ is also used across our experiments. Both the restoration and diffusion networks are trained on $128\times 128$ patches, which are processed with random cropping and rotation for data augmentation. Our model is trained with a fixed learning rate $-55e\text{-}$ 5 using Adam (Kingma & Ba, 2014) algorithm and the batch size is set to 40.
Metrics. The performance of various methods is mainly evaluated using the classical metrics: PSNR, SSIM, and LPIPS. For the image deraining, we calculate PSNR/SSIM using the Y channel in YCbCr color space following existing methods (Jiang et al., 2020; Purohit et al., 2021; Zamir et al., 2022).
4.1 Comparisons with State-of-the-Art Methods
We implement the image restoration network using a handy and classical U-Net architecture, which is trained with the proposed noise-space domain adaptation strategy. To validate its effectiveness, we compare the proposed method with previous domain adaptation approaches, including DANN (Ganin & Lempitsky, 2015), DSN (Bousmalis et al., 2016), PixelDA (Bousmalis et al., 2017), and CyCADA (Hoffman et al., 2018), covering the feature-space and pixel-space solutions. For the purpose of a fair comparison, we retrained these methods with the same standard settings and datasets. Besides, we also consider some unsupervised restoration methods and representative supervised methods such as Ne2Ne (Huang et al., 2021), MaskedD (Chen et al., 2023), NLCL (Ye et al., 2022), SelfDeblur (Ren et al., 2020), VDIP (Huo et al., 2023), and Restormer (Zamir et al., 2022).
Table 2: Quantitative evaluation of the image deraining task on SPA test dataset (Wang et al., 2019).
Metrics Vanilla DANN DSN PixelDA CyCADA NLCL Restormer Ours
@@ 4.1 Comparisons with State-of-the-Art Methods
Table 3: Quantitative evaluation of the image deblurring task on RealBlur-J test dataset (Rim et al., 2020).
Metrics Vanilla DANN DSN PixelDA CyCADA SelfDeblur VDIP Ours
@@
Figure 5: Visual comparison of the image deraining and image deblurring tasks on SPA Wang et al. (2019) and RealBlur-J Rim et al. (2020) test datasets.
Comparison Results. The quantitative and qualitative comparison results are shown in Tab. 1-3 and Fig. 4-5. From the comparison results, the proposed method leads the comparison methods on three image restoration tasks. In particular, previous feature-space domain adaptation methods (Ganin & Lempitsky, 2015; Bousmalis et al., 2016; Hoffman et al., 2018) fail to perceive the crucial low-level information and pixel-space domain adaptation methods (Bousmalis et al., 2017; Hoffman et al., 2018) yield inferior results since the precise style transfer between two domains is hard to control during the adversarial training. Moreover, the self-supervised and unsupervised restoration methods (Huang et al., 2021; Chen et al., 2023; Ye et al., 2022; Huo et al., 2023) show noticeable artifacts and limited generalization performance due to some inevitable information loss and hand-crafted designs on specific degradations. By contrast, our method ensures a general and fine domain adaptation in the pixel-wise noise space across various tasks, without introducing unstable training.
@@
Analysis. From the above results, we can observe that the proposed method enables noticeable improvements beyond the Vanilla baseline (trained only with synthetic datasets) on the tasks involved with high-frequency noises, such as image denoising. In particular, $+8.13/0.3070$ improvements on PSNR/SSIM metrics are achieved. We argue that the target of image denoising naturally fits that of the forward denoising process in the diffusion model. It is more sensitive to other Gaussian-like noises with respect to the pre-sampled noise space. Thus, an intense diffusion loss would be back-propagated if the conditioned images are under-restored, and the preceding restoration network tries to eliminate the noises on both the synthetic and real-world images as much as possible.
4.2 Ablation Studies
To evaluate the effectiveness of different components in the proposed method, we conduct ablation studies regarding the sampled noise levels of the diffusion model, determined by the time-step $t$, and the training strategies to avoid shortcut learning, as shown in Tab. 4 and Fig. 6. Concretely, with low noise intensity, e.g., $t\in[1,100]$, it is easy for the diffusion model to discriminate the similarity of paired synthetic data even when the restored conditions are under-restored. As a result, the shortcut learning occurs earlier during the training process and the real-world degraded image is heavily corrupted by the restoration network, of which most all details are filtered. On the other hand, when the intensity of the sampled noise is high, e.g., $t\in[900,1000]$, the diffusion model is hard to converge and the whole framework has fallen into a local optimum. By sampling the noise from a more diverse range with $t\in[1,1000]$, the restored results can be gradually adapted to the target clean distribution. Moreover, the generalization ability of the restoration network gains further improvement using the designed channel shuffling layer (CS) and residual-swapping contrastive learning strategy (RS), which effectively eliminates the shortcut learning of the diffusion model. Therefore, higher restoration performance on real-world images and more realistic visual appearance can be observed from (d) to (e) and (f) in Tab. 4 and Fig. 6. We also demonstrate that both synthetic data and real data are indispensable for domain adaptation in diffusion, excluding each of them would lead to dramatic degradation in real-world performance (shown in the last two rows in Tab. 4). Particularly for excluding the real data, the performance is almost degraded to that of the Vanilla model. More analysis can be found in the Appendix A3.
@@ 4.2 Ablation Studies
Figure 6: Visual comparison results of ablation studies. The complete version (f) of the proposed method achieves the best restoration results with visually pleasant appearances.
Metrics C2N AP-BSN Ours Ours*
Space - - Noise Noise
Type SS SS DA DA
Train Data both real both both
PSNR $\uparrow$ 35.35 34.90 34.71 35.52
SSIM $\uparrow$ 0.9370 0.9000 0.9202 0.9297
Table 5: Ours* denotes using a more advanced restoration network with deeper layers, trained by our domain adaptation strategy. SS and DA represent the self-supervised and domain adaptation methods, respectively. The asymmetric pixel-shuffle downsampling for the blind-spot network is exploited.
@@ 4.3 Scalability
Comparisons. In this work, we aim to present a general domain adaptation strategy for various image restoration tasks, which is scalable to any restoration network. In particular, a basic and lightweight U-Net is used to validate the effectiveness of the proposed strategy. However, such an architecture essentially limits the upper bound of the restoration performance compared to some recent promising self-supervised works (Jang et al., 2021; Lee et al., 2022; Jang et al., 2024) tailored to specific tasks.
Here, we provide experiments to demonstrate higher performance can be achieved using advanced restoration networks with the proposed adaptation strategy. The comparison results are shown in Table 5. In this experiment, we employ a restoration network based on U-Net architecture with deeper layers (named Ours*, the complexity details of different restoration networks are listed in Table A2 of Appendix). The results demonstrate that denoising performance on the SIDD test set has been improved from 34.71 dB to 35.52 dB. Moreover, we show the proposed method can generalize well to other unseen real-world datasets in Fig. 8. These datasets are not encountered during the network’s training and fall outside the distribution of the trained datasets.
We believe more powerful restoration networks can enable further improvements, but pursuing extraordinary performance for specific tasks is not the goal of this work.
Discussion. Compared to the self-supervised methods (Chen et al., 2023; Ren et al., 2020; Huo et al., 2023; Jang et al., 2021; Lee et al., 2022), our work shows the following unique strengths: it is not bounded to the specific tasks; it is free to the prior knowledge of underlying noise distribution and degradation mode; and it is friendly to the type of preceding restoration networks.
We also argue the difference between domain adaptation and self-supervised learning methods: Domain adaptation transfers knowledge from one domain to another with different data distributions, improving performance in new, unseen environments. Self-supervised learning, on the other hand, learns from unlabeled data by generating pseudo-labels or exploring the target distribution from the data itself. Both approaches reduce the reliance on large amounts of labeled data but address different challenges: domain adaptation focuses on bridging domain gaps, and self-supervised learning leverages data’s inherent structure.
@@
Peformance vs. Complexity. We then validate the scalability of the proposed method using different variants of U-Net-based image restoration networks and other types of architectures, such as the Transformer-based network (Wang et al., 2022). In particular, we classify these networks based on their model sizes and obtain: Unet-T, Unet-S (the model applied in Sec. 4.1), Unet-B, Uformer-T, Uformer-S, and Uformer-B. More network details are listed in the Appendix A1.3. The quantitative results of PSNR vs. computational costs are shown in Fig. 7. As we can observe, as the complexity and parameter increase, the vanilla restoration network (orange elements) tends to overfit the training synthetic dataset and perform worse on the test real-world dataset. In contrast, the proposed method can improve the generalization ability of restoration models with various sizes (blue elements). It is also interesting that for each type of architecture, our method can facilitate better performance as the complexity of the restoration network increases, demonstrating its effectiveness in addressing the overfitting problem of large models.
Figure 8: Visual results of the proposed method on unseen real-world datasets: the denoising test dataset DND (Plotz & Roth, 2017) and deraining test dataset ‘Real-Internet’ (Yang et al., 2017).
4.4 Limitation
The natural mission of the diffusion model is to predict the noises mixed in the input, which are sampled from a high-frequency distribution. Diffusion models excel at capturing and modeling these small-scale variations due to their ability to learn fine-grained details through their denoising process. Thus, more intuitive improvements can be observed in image denoising and deraining tasks, which typically involve high-frequency noises in images. By contrast, artifacts in blurred images, which consist of smooth, gradual changes in intensity, can be less sensitive for diffusion models. They affect larger regions of the image and require the model to correct broad, sweeping distortions rather than fine details. Consequently, diffusion models may struggle to fully restore images with low-frequency noise compared to those with high-frequency noise. We leave it as one of the future work.